# Projected state ensemble of a generic model of many-body quantum chaos

### v1

@@ VIII Conclusions
This work investigates the problem of deep thermalization in the context of random unitary circuits (RUC), by mapping the frame potential to a statistical mechanical problem. We connect the calculation of the frame potential with the moments of state purity using the replica trick, revealing the emergence of a membrane as a key feature. Through our analysis, we estimate the time scales for the emergence of deep thermalization for the $k$-th moment to be $t_{k}=O(\log k)$, a prediction that seems to be robust for generic chaotic quantum circuit as long as the membrane picture is applicable. For outlooks, it would be of interest to understand the behaviour of deep thermalization in the presence of symmetries and conserved quantities Rakovszky et al. (2018); Khemani et al. (2017); Friedman et al. (2019), to further understand the (subleading-in-$q$) differences between deep thermalization and state purity by refining the replica analysis, and to obtain an analytical approach of deep thermalization in RUC at finite-$q$ using the non-Hermitian Ginibre ensemble via space-time duality Shivam et al. (2023).
@@ Appendix A Projected ensemble for a Haar-random state
Consider the situation where the system has been prepared in a Haar-distributed random state |$\ket{\Psi}$. As explained in the main text, we use the replica trick to write the generalized frame potential in terms of $2(n+k)$ replicas and consider, in the end, the limit $n\to 1-k$. Here, we work on an arbitrary $q$. The Haar average over |$\ket{\Psi}$ gives
|$\overline{\ket{\Psi}\bra{\Psi}^{\otimes(2n+2k)}}=\frac{1}{\mathcal{N}_{2n+2k}(q^{L})}\sum_{\sigma\in\mathcal{S}_{2n+2k}}\ket{\sigma}^{\otimes L}$ (56)
Now, we use the boundary conditions on the top, which lead to
$F^{(n,k)}=\frac{1}{\mathcal{N}_{2n+2k}(q^{L})}\sum_{\sigma\in\mathcal{S}_{2n+2k}}\braket{A}{\sigma}^{L_{A}}\braket{B}{\sigma}^{L_{B}}$ (57)
@@
$\mathcal{N}_{m}(q)=q(q+1)(q+2)\ldots(q+m-1)$ (58)
@@
$\braket{B}{\sigma}=\sum_{a,a^{\prime}}\mathbf{1}_{F}(\sigma)+(1-\mathbf{1}_{F}(\sigma))\delta_{a,a^{\prime}}=\mathbf{1}_{F}(\sigma)(q^{2}-q)+q$ (59)
where $\mathbf{1}_{F}(\sigma)=1$ if $\sigma$ is factorized as $\sigma_{1}\times\sigma_{2}$, with $\sigma_{1,2}\in\mathcal{S}_{n+k}$ and 00 otherwise. We also have
$\braket{B}{\sigma}^{L_{B}}=(q^{2L_{B}}-q^{L_{B}})\mathbf{1}_{F}(\sigma)+q^{L_{B}}$ (60)
@@
$\sum_{\sigma\in\mathcal{S}_{2n+2k}}\braket{A}{\sigma}^{L_{A}}=\sum_{\sigma\in\mathcal{S}_{2n+2k}}q^{L_{A}|\mu_{A}\sigma^{-1}|}=\mathcal{N}_{2n+2k}(q^{L_{A}})$ (61)
Then, we have the second contribution
$\sum_{\sigma\in\mathcal{S}_{2n+2k}}\braket{A}{\sigma}^{L_{A}}\mathbf{1}_{F}(\sigma)=\sum_{\sigma_{1},\sigma_{2}\in\mathcal{S}_{n+k}}\braket{A}{\sigma_{1},\sigma_{2}}^{L_{A}}$ (62)
To evaluate this last term, we notice that the permutation in $A$ acts as the identity over the first/last $n$ copies and it acts as $\imath_{2k}$ on the inner $2k$ copies. Thus, we can reduce the sum over $\mathcal{S}_{n+k}$ to one in $\mathcal{S}_{k}$ using that
$\sum_{\sigma\in\mathcal{S}_{n+k}}\braket{1_{n}}{\otimes\mathbf{1}_{k}}{\sigma}^{L_{A}}=Z\sum_{\tilde{\sigma}\in\mathcal{S}_{k}}\ket{\tilde{\sigma}}^{\otimes L_{A}}$ (63)
The multiplicative constant $Z$ can be fixed by contracting both side over the identity on the residual $k$ copies. One has therefore
$\mathcal{N}_{n+k}(q^{L_{A}})=\sum_{\sigma\in\mathcal{S}_{n+k}}q^{L_{A}|\sigma|}=Z\sum_{\tilde{\sigma}\in\mathcal{S}_{k}}q^{L_{A}|\tilde{\sigma}|}=Z\mathcal{N}_{k}(q^{L_{A}})\quad\Rightarrow\quad Z=\frac{\mathcal{N}_{n+k}(q^{L_{A}})}{\mathcal{N}_{k}(q^{L_{A}})}$ (64)
So, we have
$\sum_{\sigma\in\mathcal{S}_{2n+2k}}\braket{A}{\sigma}^{L_{A}}\mathbf{1}_{F}(\sigma)=$ $\left(\frac{\mathcal{N}_{n+k}(q^{L_{A}})}{\mathcal{N}_{k}(q^{L_{A}})}\right)^{2}\sum_{\tilde{\sigma}_{1},\tilde{\sigma}_{2}\in\mathcal{S}_{k}}\braket{\imath_{2k}}{\tilde{\sigma}_{1},\tilde{\sigma}_{2}}^{L_{A}}$ (65)
$=$ $\left(\frac{\mathcal{N}_{n+k}(q^{L_{A}})}{\mathcal{N}_{k}(q^{L_{A}})}\right)^{2}\sum_{\tilde{\sigma}_{1},\tilde{\sigma}_{2}\in\mathcal{S}_{k}}q^{L_{A}|\tilde{\sigma}_{1}\tilde{\sigma}_{2}^{-1}|}=k!\frac{\mathcal{N}_{n+k}(q^{L_{A}})^{2}}{\mathcal{N}_{k}(q^{L_{A}})}$
@@
$F^{(n,k)}=\frac{1}{\mathcal{N}_{2n+2k}(q^{L})}\left(q^{L_{B}}\mathcal{N}_{2n+2k}(q^{L_{A}})+(q^{2L_{B}}-q^{L_{B}})k!\frac{\mathcal{N}_{n+k}(q^{L_{A}})^{2}}{\mathcal{N}_{k}(q^{L_{A}})}\right)$ (66)
@@
$F^{(k)}=\lim_{n\to 1-k}F^{(n,k)}=\frac{1}{\mathcal{N}_{2}(q^{L})}\left(q^{L_{B}}\mathcal{N}_{2}(q^{L_{A}})+(q^{2L_{B}}-q^{L_{B}})k!\frac{\mathcal{N}_{1}(q^{L_{A}})^{2}}{\mathcal{N}_{k}(q^{L_{A}})}\right)=\\ =\frac{1}{q^{L}(q^{L}+1)}\left(q^{L_{B}}q^{L_{A}}(q^{L_{A}}+1)+\frac{(q^{2L_{B}}-q^{L_{B}})k!q^{2L_{A}}}{\mathcal{N}_{k}(q^{L_{A}})}\right)$ (67)
In the limit $L_{B}\to\infty$, with $L=L_{A}+L_{B}$, we have
$F^{(k)}=\frac{k!}{\mathcal{N}_{k}(q^{L_{A}})}$ (68)
If we take the limit of large $L_{B}$ before $n\to 1-k$, we have
$F^{(n,k)}\stackrel{{\scriptstyle L_{B}\to\infty}}{{=}}k!\frac{q^{2L_{B}}}{q^{L(2n+2k)}}\frac{\mathcal{N}_{n+k}(q^{L_{A}})^{2}}{\mathcal{N}_{k}(q^{L_{A}})}\stackrel{{\scriptstyle q\to\infty}}{{=}}\frac{1}{q^{L_{B}(2n+2k-2)}}\frac{k!}{q^{kL_{A}}}$ (69)

### v2

@@ VIII Conclusions
This work investigates the problem of deep thermalization in the context of random unitary circuits (RUC), by mapping the frame potential to a statistical mechanical problem. We connect the calculation of the frame potential with the moments of state purity using the replica trick, revealing the emergence of a membrane as a key feature. Through our analysis, we estimate the time scales for the emergence of deep thermalization for the $k$-th moment to be $t_{k}=O(\log k)$, a prediction that seems to be robust for generic chaotic quantum circuit as long as the membrane picture is applicable. For outlooks, it would be of interest to understand the behaviour of deep thermalization in the presence of symmetries and conserved quantities Rakovszky et al. (2018); Khemani et al. (2017); Friedman et al. (2019); Mark et al. (2024): In particular symmetry breaking can give rise to prethermalisation (or more generally slow thermalisation) Bertini et al. (2015); Kormos et al. (2017) which is expected to be seen also in the higher-replica observables studied here. It is also of interest to understand the subleading-in-$q$ differences between deep thermalization and state purity by refining the replica analysis, which can perhaps be investigated using the non-Hermitian Ginibre ensemble via space-time duality Shivam et al. (2023).
IX Acknowledgements
[…]
@@ Appendix A Projected ensemble for a Haar-random state
Consider the situation where the system has been prepared in a Haar-distributed random state |$\ket{\Psi}$. As explained in the main text, we use the replica trick to write the generalized frame potential in terms of $2(n+k)$ replicas and consider, in the end, the replica limit $n\to 1-k$. Here, we work with arbitrary value of $q$. The Haar average over |$\ket{\Psi}$ gives
(|$\overline{\left(\ket{\Psi}\bra{\Psi}\right)^{\otimes(2n+2k)}}=\frac{1}{\mathcal{N}_{2n+2k}(q^{L})}\sum_{\sigma\in\mathcal{S}_{2n+2k}}\ket{\!\ket{\sigma}}^{\otimes L}\,,$ (56)
@@
$\mathcal{N}_{m}(q):=q(q+1)(q+2)\ldots(q+m-1)\,.$ (57)
By using Eq. (56) and a straightforward generalization of (38), and by applying the boundary conditions at the top in Fig. 2, we obtain
$F^{(n,k)}=\frac{1}{\mathcal{N}_{2n+2k}(q^{L})}\sum_{\sigma\in\mathcal{S}_{2n+2k}}\braket{\!\braket{A}{\sigma}}{}{L_{A}}\braket{\!\braket{B}{\sigma}}{}{L_{B}}\,,$ (58)
@@
$\braket{\!\braket{B}{\sigma}}{=}\sum_{a,a^{\prime}}\mathbf{1}_{F}(\sigma)+(1-\mathbf{1}_{F}(\sigma))\delta_{a,a^{\prime}}=\mathbf{1}_{F}(\sigma)(q^{2}-q)+q\,.$ (59)
where $\mathbf{1}_{F}(\sigma)=1$ if $\sigma$ is factorized as $(\sigma_{1},\sigma_{2})$, with $\sigma_{1},\sigma_{2}\in\mathcal{S}_{n+k}$ and 00 otherwise. We also have
$\braket{\!\braket{B}{\sigma}}{}{L_{B}}=(q^{2L_{B}}-q^{L_{B}})\mathbf{1}_{F}(\sigma)+q^{L_{B}}\,,$ (60)
@@
$\sum_{\sigma\in\mathcal{S}_{2n+2k}}\braket{\!\braket{A}{\sigma}}{}{L_{A}}=\sum_{\sigma\in\mathcal{S}_{2n+2k}}q^{L_{A}N_{c}(\mu_{A}\sigma^{-1})}=\mathcal{N}_{2n+2k}(q^{L_{A}})\,,$ (61)
where, as in the main text, $N_{c}(\sigma)$ counts the number of cycles in the permutation $\sigma\in S_{m}$. The second contribution is
$\sum_{\sigma\in\mathcal{S}_{2n+2k}}\braket{\!\braket{A}{\sigma}}{}{L_{A}}\mathbf{1}_{F}(\sigma)=\sum_{\sigma_{1},\sigma_{2}\in\mathcal{S}_{n+k}}\braket{\!\braket{A}{(\sigma_{1},\sigma_{2})}}{}{L_{A}}\,.$ (62)
To evaluate this last term, we notice that the permutation in $A$ acts as the identity over the first/last $n$ copies and it acts as $\imath_{2k}$ on the inner $2k$ copies (see Fig. 2). We can use this fact to effectively reduce the sums in $\mathcal{S}_{n+k}$ to sums in $\mathcal{S}_{k}$. To do so, we denote as $\left(\bra{\!\bra{1_{n}}}\otimes\mathbb{1}_{k}\right)$ the projector on the identity permutation for the first $n$ replicas, where $\mathbb{1}_{k}$ denotes the identity operator on the remaining $k$ replicas. Then, we have the identity
[…]
This equality can be simply understood noting that the left-hand side is invariant under any permatution of the last $k$ replicas and so it must be proportional to the flat sum over all elements in the group appearing in the right-hand side. The multiplicative constant $Z$ can be fixed by contracting both side over the identity on the residual $k$ copies,
$\mathcal{N}_{n+k}(q^{L_{A}})=\sum_{\sigma\in\mathcal{S}_{n+k}}q^{L_{A}N_{c}(\sigma)}=Z\sum_{\tilde{\sigma}\in\mathcal{S}_{k}}q^{L_{A}N_{c}(\tilde{\sigma})}=Z\mathcal{N}_{k}(q^{L_{A}})\quad\Rightarrow\quad Z=\frac{\mathcal{N}_{n+k}(q^{L_{A}})}{\mathcal{N}_{k}(q^{L_{A}})}\,.$ (64)
Substituting this back into (63), we have
$\sum_{\sigma\in\mathcal{S}_{2n+2k}}\braket{\!\braket{A}{\sigma}}{}{L_{A}}\mathbf{1}_{F}(\sigma)=$ $\left(\frac{\mathcal{N}_{n+k}(q^{L_{A}})}{\mathcal{N}_{k}(q^{L_{A}})}\right)^{2}\sum_{\tilde{\sigma}_{1},\tilde{\sigma}_{2}\in\mathcal{S}_{k}}\braket{\!\braket{\imath_{2k}}{(\tilde{\sigma}_{1},\tilde{\sigma}_{2})}}{}{L_{A}}$ (65)
$=$ $\left(\frac{\mathcal{N}_{n+k}(q^{L_{A}})}{\mathcal{N}_{k}(q^{L_{A}})}\right)^{2}\sum_{\tilde{\sigma}_{1},\tilde{\sigma}_{2}\in\mathcal{S}_{k}}q^{L_{A}N_{c}(\tilde{\sigma}_{1}\tilde{\sigma}_{2}^{-1})}=k!\frac{\mathcal{N}_{n+k}(q^{L_{A}})^{2}}{\mathcal{N}_{k}(q^{L_{A}})}\,.$
@@
$F^{(n,k)}=\frac{1}{\mathcal{N}_{2n+2k}(q^{L})}\left[q^{L_{B}}\mathcal{N}_{2n+2k}(q^{L_{A}})+(q^{2L_{B}}-q^{L_{B}})k!\frac{\mathcal{N}_{n+k}(q^{L_{A}})^{2}}{\mathcal{N}_{k}(q^{L_{A}})}\right]\,.$ (66)
@@
$F^{(k)}=\lim_{n\to 1-k}F^{(n,k)}=\frac{1}{\mathcal{N}_{2}(q^{L})}\left[(q^{L_{B}}\mathcal{N}_{2}(q^{L_{A}})+(q^{2L_{B}}-q^{L_{B}})k!\frac{\mathcal{N}_{1}(q^{L_{A}})^{2}}{\mathcal{N}_{k}(q^{L_{A}})}\right]=\\ =\frac{1}{q^{L}(q^{L}+1)}\left[(q^{L_{B}}q^{L_{A}}(q^{L_{A}}+1)+\frac{(q^{2L_{B}}-q^{L_{B}})k!q^{2L_{A}}}{\mathcal{N}_{k}(q^{L_{A}})}\right]\,.$ (67)
With $L=L_{A}+L_{B}$, in the limit $L_{B}\to\infty$, we have
$F^{(k)}=\frac{k!}{\mathcal{N}_{k}(q^{L_{A}})}\,.$ (68)
Further, we can verify that the limits involved in $L_{B}\to\infty,n\to 1-k$ and our approximation $q\to\infty$ commutes among themselves,
$F^{(n,k)}\stackrel{{\scriptstyle L_{B}\to\infty}}{{=}}k!\frac{q^{2L_{B}}}{q^{L(2n+2k)}}\frac{\mathcal{N}_{n+k}(q^{L_{A}})^{2}}{\mathcal{N}_{k}(q^{L_{A}})}\stackrel{{\scriptstyle q\to\infty}}{{=}}\frac{1}{q^{L_{B}(2n+2k-2)}}\frac{k!}{q^{kL_{A}}}\stackrel{{\scriptstyle n\to 1-k}}{{=}}\frac{k!}{q^{kL_{A}}}\,.$ (69)
Note that the large-$q$ expansion is formally equivalent to the large–$L_{A}$ expansion for a Haar-random state. However, for generic local quantum circuits, the time-dependence of the frame potential explicitly depends on $L_{A}$ even though $q\to\infty$.